OpenStudy (anonymous):

What is the value of f(x) = 9x when x = -2?

OpenStudy (catlover5925):

9 * -2 = -18

OpenStudy (anonymous):

are you sure its not positive ? /:@Catlover5925

OpenStudy (anonymous):

@Catlover5925

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x)=9x }\) to find f(-3), you plug in -3 for x. \(\large\color{black}{ f(-3)=9(-3)=-27 }\) so \(\large\color{black}{ f(-3)=-27 }\) to find f(1), you plug in 1 for x. \(\large\color{black}{ f(1)=9(1)=9 }\) so \(\large\color{black}{ f(1)=9 }\) to find f(0), you plug in 0 for x. \(\large\color{black}{ f(0)=9(0)=0 }\) so \(\large\color{black}{ f(0)=0 }\) see? to find f(-2), you plug in -2for x. \(\large\color{black}{ f(-2)=9(-2)=? }\) so \(\large\color{black}{ f(-2)=? }\)

OpenStudy (anonymous):

ohh okkay thank @SolomonZelman so it would be -18? but my answer choices dont have a negative 18 only positive 18

OpenStudy (solomonzelman):

yes it is -18.

OpenStudy (catlover5925):

why do you think it would be positive??? it is a negative * a positive.... if it was a negative * negative then it would be positive

OpenStudy (solomonzelman):

maybe it is \(\large\color{black}{ f(x)=9^x }\) ?

OpenStudy (anonymous):

yes that @SolomonZelman

OpenStudy (solomonzelman):

so to find f(-2), plug in -2 for x into the function.

OpenStudy (solomonzelman):

There is a rule, \(\LARGE\color{blue}{ a^{-c}= \frac{1}{a^c} }\) and applying this rule to \(\Large\color{black}{ 9^{-2}= \frac{1}{a} }\) what do you get?

OpenStudy (solomonzelman):

\(\Large\color{black}{ 9^{-2}= ? }\)

OpenStudy (catlover5925):

oh ok now that makes more sence!! now it would be positive

OpenStudy (solomonzelman):

yes:) but an 18 though:)

OpenStudy (anonymous):

alright thanks guys @Catlover5925 @SolomonZelman